Find the Wronskian of the following set of function
{1-x,1+x, 1-3x}
1
Expert's answer
2021-09-21T12:23:49-0400

f1=1-x

f2=1+x

f3=1-3x

"W=\begin{bmatrix}
 f_1 & f_2 & f_3 \\
 f_1' & f_2' & f_3' \\
f_1'' & f_2'' & f_3''
\end{bmatrix}"

"W=\begin{bmatrix}
 1-x & 1+x & 1-3x \\
 -1 & 1 & -3 \\
0 & 0 & 0
\end{bmatrix}"

=(1-x)"\begin{bmatrix}
 1 & -3 \\
 0 & 0
\end{bmatrix}" -(1+x)"\begin{bmatrix}
 -1 & -3 \\
 0 & 0
\end{bmatrix}" +(1-3x)"\begin{bmatrix}
 -1 & 1 \\
 0 & 0
\end{bmatrix}"

=(1-x)(0-0)-(1-x)(0-0)+(1-3x)(0-0)

=0